# CaloQVAE : Simulating high-energy particle-calorimeter interactions using hybrid quantum-classical generative models

### v3

@@ 1 Introduction
Recently, significant advancements have been made in various machine-learning domains related to generative models. In the development of variational autoencoders (VAEs) and their extensions, a series of breakthroughs have shaped the field kingma2022autoencoding ; rolfe2016discrete ; vahdat2018dvae++ ; khoshaman2018gumbolt . The development of the generative models provides inspiration to collider physics in the context of calorimeter simulation. In 2018, novel techniques based on generative adversarial networks (GANs) were developed to address the need for fast simulation of electromagnetic showers in calorimeters CaloGAN . In 2021, the advancements have introduced innovative frameworks like CaloFlow based on normalizing flows, offering fast and accurate simulations with high fidelity and stability CaloFlow . Most recently, IEA-GAN combines Self-Supervised Learning with Generative Adversarial Networks to efficiently simulate ultra-high-resolution particle detector responses with a relational reasoning module IEA-GAN .
@@
In traditional generative models, the computational costs and the time required for GPU processing in Markov Chain evaluations are significant concerns. Therefore, we aim to develop a quantum-assisted generative model for calorimeter data simulation, inspired by the potential of quantum computing to reduce these computational burdens. In 2015, a method utilising quantum annealing for training deep neural networks was proposed with a description on how to map Restricted Boltzmann Machines onto a D-Wave Two hardware Adachi2015 . Building on this idea, a subsequent work explored quantum annealers’ ability to sample from Boltzmann distributions effective_temperatures_estimation . The development of the quantum variational autoencoder (QVAE) showcased the feasibility of utilising a quantum Boltzmann machine as a prior for latent generative process in a variational autoencoder Khoshaman_2018 ; Winci_2020 . In 2021, a hybrid quantum-classical qGAN showed promise for accelerating HEP detector simulations using quantum generator circuits and classical discriminator neural networks qGAN . In a very recent development, CaloDVAE introduced a discrete variational autoencoder with an RBM prior for fast calorimeter shower simulation, yielding promising results in generating realistic and diverse samples. The possibility of using quantum annealing processors as sampling devices for the RBM prior was also discussed Abhi .
@@
In this paper, we extend the previous work from CaloDVAE to CaloQVAE by employing the D-Wave quantum hardware DWAVE to get samples from the latent space in a classically trained Variational Autoencoder (VAE), thereby bypassing the need for computationally intensive Markov Chain evaluations used in shower model simulations. To our knowledge, this approach marks the first instance of utilizing quantum computing to generative models for calorimeter shower simulations at a realistic scale.
@@ 2 Model Setup
CaloDVAE rolfe2016discrete is a hierarchical discrete VAE which extends the VAE framework by introducing discrete latent variables $\mathbf{z}_{i}\in\{0,1\}$ in the latent space. Equation 1 is the loss for the unconditional VAE model. For conditional VAEs the encoder and decoder have an additional conditioning variable. In our case the encoder, $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{e})$, and decoder, $p_{\theta}(\mathbf{x}|\mathbf{z},\mathbf{e})$, are conditioned on energy $\mathbf{e}$ Abhi ; energy-conditioning (unlike the encoders and decoders of Equation 1 of the unconditional VAE framework) and are modelled by fully connected neural networks. The approximate posterior of the VAE has a hierarchical structure $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{e})=\prod_{i}q_{\phi_{i}}(\mathbf{z}_{i}|\mathbf{z}_{j<i},\mathbf{x},\mathbf{e})$ with $N$ latent groups and the latent distribution of the VAE is modelled by an RBM where $\mathbf{z}=[\mathbf{z}_{1},...,\mathbf{z}_{N}]$ are partitioned into $2$ equal subsets to make the $2$ sides of the RBM Abhi . The RBM is modeled by probability distribution $p_{RBM}=e^{\mathbf{a}_{l}^{T}\mathbf{z}_{l}+\mathbf{a}_{r}^{T}\mathbf{z}_{r}+\mathbf{z}_{l}^{T}\mathbf{W}\mathbf{z}_{r}}/Z$ where $Z$ is the partition function and $(\mathbf{a}_{l},\mathbf{a}_{r},\mathbf{W})$ are the RBM parameters which are trained along with the VAE parameters $(\theta,\phi)$.
@@ 3 CaloQVAE Training and Inference
In this work we demonstrate that we can generate latent space samples using the D-Wave 2000Q annealer. The first step to use the quantum processing unit (QPU) in the pipeline is to embed the RBM on the quantum hardware. However, the QPU architecture is not fully connected and instead follows a Chimera graph topology, making it impractical to natively embed a fully connected RBM in the QPU Adachi2015 . We therefore create a masking function to remove selected edges from an RBM to produce a Chimera RBM. We train the model classically analogously to the CaloDVAE training by replacing the RBM of CaloDVAE with a Chimera RBM.
@@
In the inference phase we first convert the Chimera RBM Hamiltonian to an Ising model Hamiltonian $\mathcal{H}_{P}$ using a linear transformation Dixit2022 . We embed the Chimera RBM on the QPU, and anneal according to the QPU Hamiltonian $\mathcal{H}(s)=A(s)\mathcal{H}_{T}+B(s)\mathcal{H}_{P}$, where $\mathcal{H}(s)$ is the total Hamiltonian of system at the re-scaled time $s$ of annealing, $\mathcal{H}_{T}$ is the transverse field Hamiltonian, $A(s)$ and $B(s)$ are monotonic functions amin2018quantum . We anneal $\mathcal{H}(s)$ for $20$ $\mathrm{\mu s}$ following the standard D-Wave annealing schedule. During the anneal a freezing point $s^{*}$ is typically observed on the device AminFreeze ; Marshall2017 , after which $A(s^{*})\ll B(s^{*})$ and the change in dynamics becomes negligible. If the time taken to freeze is sufficiently short, the device output configurations $x$ will approximately follow a Boltzmann Distribution at $s^{*}$, with $P(x)\propto\exp{(-\beta_{eff}^{*}\mathcal{H}_{P}(x))}$. This defines an effective inverse temperature $\beta_{eff}^{*}$ Adachi2015 ; Perdomo2015 which is discovered to be instance-dependent and fluctuate at different programming cycles freeze_out . Therefore, a real-time $\beta_{eff}^{*}$ estimation is needed XuAdaptive2021 . In our work we estimate $\beta_{eff}^{*}$ using an iterative procedure Winci_2020 where we first generate samples from the trained Chimera RBM classically using Block Gibbs Sampling. We then iteratively generate synthetic samples from the QPU while adjusting $\beta_{eff}^{*}$ and then rescaling the QPU parameters (couplings and local fields) by $\beta_{eff}^{*}$ until the quantum energy distribution converges to the classical energy distribution Winci_2020 . Figure 1 shows energy distributions for the QPU samples and classical samples.
@@ 4 Performance Evaluation
Figure 2: Comparison of shower shape variables for $\pi^{+}$ and $e^{+}$ particles in a calorimeter. This figure shows the distributions of two key variables: (1) Fraction of energy: The energy deposited in the first layer compared to the sum of energy in all layers. (2) Sparsity: The ratio of the number of hits in the first layer to the total number of hits in all layers. Each variable is plotted for both particle types to illustrate the close alignment among the results from GEANT4, DVAE (classical) and QVAE (quantum) models.
@@
For training and performance evaluation we use the public CaloGAN dataset Nachman2017 simulated from GEANT4 GEANT4:2002zbu . The dataset provides 100,000 events for $\pi^{+}$, $\gamma$ and $e^{+}$ incident particles CaloGAN and is split into 80%, 10%, and 10% subsets for training, validation, and testing, respectively. After conducting a hyperparameter scan Abhi , we obtain the best models for each particle type. Utilizing classical blocked Gibbs sampling and QPU sampling methods respectively, we generate new samples from the latent space and pass them through the decoder to obtain synthetic shower images. These images are then compared with GEANT4 test data to assess their performance. Classification can be used as a useful indicator of the performance of generative models. A good classifier should be able to understand and capture the important patterns in the data because it needs to learn discriminating features that distinguish different classes or categories accurately. We perform binary classification of $e^{+}$ vs $\pi^{+}$ and $e^{+}$ vs $\gamma$ using a neural network with 6 fully connected layers with dropout, following the method of CaloGAN . All synthetic data is generated from the QPU using separate models trained separately on $e^{+},\pi^{+}$ and $\gamma$ samples. Our accuracy metrics (Table 1) are very close to the results obtained in CaloGAN . The similar accuracy metrics illustrate that our CaloQVAE model is generating synthetic data that captures important patterns present in the true data, which indicates that the synthetic images generated from the QPU are successful at representing the underlying structure and characteristics of the true data. The fact that the classifier trained on synthetic $e^{+}$ vs $\pi^{+}$ data can perform well when tested on the GEANT4 data, and vice versa, suggests that the features learned by the CaloQVAE model are transferable. This means that the CaloQVAE model has successfully encoded the relevant information from the true data into its latent space representation, which allows the generated data to be informative and useful for training a classifier. An issue that requires attention is the unexpected high performance of the $e^{+}$ vs $\gamma$ classifier, which is trained with synthetic data and tested on a separate set of synthetic data. Similar over-performance has also been observed in CaloGAN . This raises the possibility that our CaloQVAE model may generate certain non-physical features, which allow the classifier to distinguish between the $e^{+}$ and $\gamma$ samples. Further research is warranted to explore potential ways of addressing this challenge.
@@
Table 1: Binary classification accuracy for $e^{+}$ vs $\pi^{+}$ and $e^{+}$ vs $\gamma$ where we train and test both on the Geant4 data and synthetic data produced from QPU samples
@@
Train Geant4 99.8 99.9 66.8 74.1
Synthetic 94.3 100.0 53.7 99.9
@@
We qualitatively analyze one-dimensional histograms of the typical shower shape variables. As shown in Fig 2, there is a close alignment between the results from GEANT4 data and generative models, which signifies that our generative model effectively captures the underlying patterns and distributions of the data. The consistent distribution patterns highlight the generative model’s fidelity in representing the essential features in the training dataset. Despite the observation of tail differences at the low fraction region for the positron in layer 1’s fraction histogram, the overall similarity between the performance of QVAE and DVAE models indicates our pipeline for Beta estimation and scaling is well behaved so that the quality of samples from the quantum device is on par with those from the Blocked Gibbs Method.
@@
To assess the practical deployment performance of the model’s energy conditioning, we generate samples from the trained model requesting a specific range of incident energy and histogram the observed energy. We generate the samples by classical and quantum approaches and compare the results with a GEANT4 test dataset under the same range of incident energy. As shown in Fig 3, for the positron cluster, the conditioning response of DVAE or QVAE can form a flat plateau with good accuracy in reproducing the defined selection of the input energy range. The pion cluster has a broadened response that is due to the nature of the unconstrained charged pion shower in the electromagnetic calorimeter. The histograms of the generative models for both electron and pion clusters are able to mirror the characteristics observed in the true data. This compelling agreement underscores the VAE’s ability to capture and faithfully reproduce the intricate patterns inherent in the first-principles simulation, thereby affirming its efficacy as a powerful tool for energy pattern generation.
@@
Figure 3: Comparison of energy spectra between CaloDVAE (classical) and CaloQVAE (quantum) models against GEANT4 test data for incident particle energies ranging from 45 to 60 GeV. The data encompasses both positron and pion clusters.
@@
The monitoring results of the DVAE sampling rate show efficient performance on our NVIDIA GeForce RTX 2080 Ti with 11264 MB memory. The processing time per sample for latent space sampling takes the most substantial portion ( $\sim$ 0.5 $\mathrm{ms}$), followed by data loading to the GPU ( $\sim$ 10 $\mathrm{ns}$), and passing through the decoder ($\sim$ 1 $\mathrm{\mu s}$). The DVAE model exhibits similar sampling rate with the CaloGAN results CaloGAN up to an order of magnitude. On the other hand, the quantum sampler achieves slightly faster sampling rate at 0.4 $\mathrm{ms}$ per sample, but impressively, the core QPU annealing rate is only 20 $\mathrm{\mu s}$ per sample, hinting at the raw speed of quantum processing. The extra QPU access time is mainly used for tasks like readout and handling delays, suggesting that with optimized engineering, quantum sampling could vastly outpace classical methods in efficiency for generative modeling, making it well-suited for real-time and large-scale applications. We expect to apply the QVAE model to more complex objects, jets and events in the future, for which annealing time is expected to remain at the current level.

### v4

@@ 1 Introduction
Recently, significant advancements have been made in various machine-learning domains related to generative models. In the development of variational autoencoders (VAEs) and their extensions, a series of breakthroughs have shaped the field kingma2022autoencoding ; rolfe2016discrete ; vahdat2018dvae++ ; khoshaman2018gumbolt . The development of the generative models provides inspiration to collider physics in the context of calorimeter simulation. In 2018, novel techniques based on generative adversarial networks (GANs) were developed to address the need for fast simulation of electromagnetic showers in calorimeters CaloGAN . Later IEA-GAN combines Self-Supervised Learning with GAN to efficiently simulate ultra-high-resolution particle detector responses with a relational reasoning module IEA-GAN . In 2021, advancements introduced innovative frameworks like CaloFlow, which are based on normalizing flows and offer fast and accurate simulations with high fidelity and stability CaloFlow . In 2022, score-based generative models were introduced for calorimeter shower simulation PhysRevD.106.092009 . In 2023, CaloFlow v2 further accelerated shower generation by a factor of 500 relative to the original by using probability density distillation PhysRevD.107.113004 . Most recently, building upon CaloFlow, L2LFlows adapted it to a higher-dimensional setting, thereby improving generative fidelity Diefenbacher_2023 .
@@
In deep generative models, the computational costs and the time required for GPU processing are significant concerns. Therefore, we aim to develop a quantum-assisted generative model for calorimeter data simulation, inspired by the potential of quantum computing to reduce these computational burdens. In 2015, a method utilising quantum annealing for training deep neural networks was proposed with a description on how to map Restricted Boltzmann Machines onto a D-Wave Two hardware Adachi2015 . Building on this idea, a subsequent work explored quantum annealers’ ability to sample from Boltzmann distributions effective_temperatures_estimation . The development of the quantum variational autoencoder (QVAE) showcased the feasibility of utilising a quantum Boltzmann machine as a prior for latent generative process in a variational autoencoder Khoshaman_2018 ; winci2020path . In 2021, a hybrid quantum-classical qGAN showed promise for accelerating HEP detector simulations using quantum generator circuits and classical discriminator neural networks qGAN . In a very recent development, CaloDVAE introduced a discrete variational autoencoder with an RBM prior for fast calorimeter shower simulation, yielding promising results in generating realistic and diverse samples. The possibility of using quantum annealing processors as sampling devices for the RBM prior was also discussed Abhi .
@@
In this paper, we extend the previous work from CaloDVAE to CaloQVAE by employing the D-Wave quantum hardware DWAVE to bypass the need for computationally intensive Markov Chain evaluations used in CaloDVAE, and get samples at a high rate from the latent space of a classically trained model. Our model enables the simulation of showers with clusters that have a granularity and number of cells, as well as energy scales, similar to those expected in actual experiments. This level of detail, corresponding to clusters observed in current and future LHC experiments, marks the first instance of utilizing quantum computing for generative models in calorimeter shower simulations at such a realistic scale.
@@ 2 Model Setup
CaloDVAE rolfe2016discrete is a hierarchical discrete VAE which extends the VAE framework by introducing discrete latent variables $\mathbf{z}_{i}\in\{0,1\}$ in the latent space. Equation 1 is the loss for the unconditional VAE model. For conditional VAEs the encoder and decoder have an additional conditioning variable. In our case the auto-regressive encoder, $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{e})$, and decoder, $p_{\theta}(\mathbf{x}|\mathbf{z},\mathbf{e})$, are conditioned on energy $\mathbf{e}$ Abhi ; energy-conditioning (unlike the encoders and decoders of Equation 1 of the unconditional VAE framework) and are modelled by fully connected neural networks. The approximate posterior of the VAE has a hierarchical structure $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{e})=\prod_{i}q_{\phi_{i}}(\mathbf{z}_{i}|\mathbf{z}_{j<i},\mathbf{x},\mathbf{e})$ with $N$ latent groups and the latent distribution of the VAE is modelled by an RBM where $\mathbf{z}=[\mathbf{z}_{1},...,\mathbf{z}_{N}]$ are partitioned into $2$ equal subsets to make the $2$ sides of the RBM Abhi . The RBM is modeled by probability distribution $p_{RBM}=e^{\mathbf{a}_{l}^{T}\mathbf{z}_{l}+\mathbf{a}_{r}^{T}\mathbf{z}_{r}+\mathbf{z}_{l}^{T}\mathbf{W}\mathbf{z}_{r}}/Z$ where $Z$ is the partition function and $(\mathbf{a}_{l},\mathbf{a}_{r},\mathbf{W})$ are the RBM parameters which are trained along with the VAE parameters $(\theta,\phi)$.
@@ 3 CaloQVAE Training and Inference
In this work we demonstrate that we can generate latent space samples using the D-Wave 2000Q annealer. Note that Quantum Annealers are mostly known to be used to find ground states of Hamiltonians to which some optimization problems can be translated to 10.1145/2482767.2482797 . However, due to interactions with the environment there is some probability that at the end of anneal, the final system will be in an excited state instead of the ground state. The probability distribution of the excited states follow the Boltzmann distribution, and hence D-Wave can also be used as an effective sampler from that distribution Adachi2015 .
@@
The first step to use the quantum processing unit (QPU) in the pipeline is to embed the RBM on the quantum hardware. However, the QPU architecture is not fully connected and instead follows a Chimera graph topology, making it impractical to natively embed a fully connected RBM in the QPU Adachi2015 . We therefore create a masking function to remove selected edges from an RBM to produce a Chimera RBM. We train the model classically analogously to the CaloDVAE training by replacing the RBM of CaloDVAE with a Chimera RBM.
@@
In the inference phase, we first convert the Chimera RBM Hamiltonian, $\mathcal{H}_{\text{P}}$, to an Ising model Hamiltonian using a linear transformation Dixit2022 . We then embed the Chimera RBM on the QPU and anneal according to the QPU Hamiltonian $\mathcal{H}(s)=A(s)\mathcal{H}_{\text{T}}+B(s)\mathcal{H}_{\text{P}}$, where $\mathcal{H}(s)$ is the total Hamiltonian of the system at the rescaled time $s$ during annealing, $\mathcal{H}_{\text{T}}$ is the transverse field Hamiltonian, and $A(s)$ and $B(s)$ are monotonic functions amin2018quantum . The annealing process follows the standard D-Wave annealing schedule and lasts for $20\ \mu\text{s}$. During this process, a freezing point $s^{*}$ is typically observed on the device AminFreeze ; Marshall2017 , after which $A(s^{*})\ll B(s^{*})$ and the system’s dynamics change negligibly. If the freezing time is sufficiently short, the resulting configurations $x$ from the device will approximately follow a Boltzmann distribution at $s^{*}$, given by $P(x)\propto\exp{(-\beta_{\text{eff}}^{*}\mathcal{H}_{\text{P}}(x))}$. This defines an effective inverse temperature $\beta_{\text{eff}}^{*}$ Adachi2015 ; Perdomo2015 , which varies with different programming cycles and depends on several unknown parameters freeze_out . Our goal is to generate samples from the distribution $P_{\text{ideal}}(x)\propto\exp{(-\mathcal{H}_{\text{P}}(x))}$, as $P_{\text{ideal}}(x)$ represents the distribution of the Chimera RBM from which we wish to sample. To achieve this, if $\beta_{\text{eff}}^{*}$ was known, the model parameters of $\mathcal{H}_{\text{P}}(x)$ could be scaled by $\beta_{\text{eff}}^{*}$ to match the desired distribution $P_{\text{ideal}}(x)$. Therefore, since $\beta_{\text{eff}}^{*}$ is not known a priori and varies with different instances, real-time estimation of $\beta_{\text{eff}}^{*}$ is necessary XuAdaptive2021 . A single run of this scaling parameter estimation is needed, after which numerous samples from the QPU can be rapidly obtained.
In our work, we estimate $\beta_{\text{eff}}^{*}$ using an iterative procedure winci2020path . The process begins with the classical generation of samples from the trained Chimera RBM using block Gibbs sampling. Then synthetic samples are iteratively generated from the QPU, and $\beta_{\text{eff}}^{*}$ is updated following an update rule winci2020path to align the mean energies of the RBM and D-Wave samples. This involves rescaling the QPU parameters (couplings and local fields) by the updated $\beta_{\text{eff}}^{*}$ for each iteration. The $\beta_{\text{eff}}^{*}$ training process is repeated over multiple ensambles, with close monitoring of the $\beta_{\text{eff}}^{*}$ values until they converge to a stable value. Figure 1 illustrates the energy distributions for both QPU and classical samples after the procedure concluded, demonstrating convergence of the quantum annealing distribution to the one obtained from the classcal RBM. In CaloQVAE, with a trained model, the $\beta_{\text{eff}}^{*}$ estimation is performed only at the start of the sampling process to determine the current temperature and does not need to be repeated for subsequent D-Wave sampling requests.
@@ 4 Performance Evaluation
Figure 2: Comparison of selected shower shape variables for positrons (left sub-panel pair) and charged pions (right sub-panel pair) in a calorimeter. Each pair of sub-panels shows the distributions of two key variables: (1) Fraction of energy: The energy deposited in the first layer compared to the sum of energy in all layers. (2) Sparsity: The ratio of the number of hits in the first layer to the total number of hits in all layers. Each variable is plotted for both particle types to illustrate the close alignment among the results from GEANT4, DVAE (classical) and QVAE (quantum) models.
@@
For training and performance evaluation we use the public CaloGAN dataset Nachman2017 simulated from GEANT4 GEANT4:2002zbu . The dataset provides 100,000 events for $\pi^{+}$, $\gamma$ and $e^{+}$ incident particles CaloGAN and is split into 80%, 10%, and 10% subsets for training, validation, and testing, respectively. After conducting a hyperparameter scan Abhi , we obtain the best models for each particle type. Utilizing classical block Gibbs sampling and QPU sampling methods respectively, we generate new samples from the latent space and pass them through the decoder to obtain synthetic shower images. These images are then compared with GEANT4 test data to assess their performance. Firstly, chi-square and KS tests were conducted on the total event energies for CaloDVAE and CaloQVAE samples in the 1-100 GeV incident energy range. The results indicated consistency between the two sets of synthetic samples, further supporting the successful utilization of the QPU for sampling from the classically trained latent space. Compared with GEANT4 samples on chi-square tests, CaloQVAE performs similarly to the ones already adopted, within an order of magnitude ATLAS:2021pzo . Moreover, classification can be used as a useful indicator of the performance of generative models. A good classifier should be able to understand and capture the important patterns in the data because it needs to learn discriminating features that distinguish different classes or categories accurately. We perform binary classification of $e^{+}$ vs $\pi^{+}$ and $e^{+}$ vs $\gamma$ using a neural network with 6 fully connected layers with dropout, following the method of CaloGAN . All synthetic data is generated from the QPU using separate models trained separately on $e^{+},\pi^{+}$ and $\gamma$ samples. Our accuracy metrics (Table 1) are very close to the results obtained in CaloGAN . The similar accuracy metrics illustrate that our CaloQVAE model is generating synthetic data that captures important patterns present in the true data, which indicates that the synthetic images generated from the QPU can represent the underlying structure and characteristics of the true data. The fact that the classifier trained on synthetic $e^{+}$ vs $\pi^{+}$ data can perform well when tested on the GEANT4 data, and vice versa, suggests that the features learned by the CaloQVAE model are transferable. This means that the CaloQVAE model can encode the relevant information from the true data into its latent space representation, which allows the generated data to be informative and useful for training a classifier. An issue that requires attention is the unexpected high performance of the $e^{+}$ vs $\gamma$ classifier, which is trained with synthetic data and tested on a separate set of synthetic data. To address the issue, we checked the datasets and found that the mean energy distributions in 504 calorimeter cells were consistent for $e^{+}$ and $\gamma$ synthetic datasets. We tried solutions like masking low-energy pixels and unifying dataset normalization, but the issue of over-performance persists. We believe that such performance difference is due to an existing intrinsic bias in the GEANT4 dataset between $e^{+}$ vs $\gamma$. As shown in Table 1, training and testing both on GEANT4 data for positron vs gamma already yield a relatively high classification rate compared to expectations. This rate increases further when training on GEANT4 and testing on synthetic data. Our model learns and amplifies this bias from the training dataset, leading to the observed issue. Similar over-performance has also been observed in CaloGAN .
@@
Table 1: Mean (standard deviation) of the binary classification accuracy over 10 particle classification trials for $e^{+}$ vs $\pi^{+}$ and $e^{+}$ vs $\gamma$ where we train and test both on the GEANT4 data and QPU synthetic data.
@@
Train Geant4 99.8 ($<$0.1) 99.8 ($<$0.1) 67.1 (0.5) 74.8 (2.8)
Synthetic 90.6 (4.0) 100.0 ($<$0.1) 53.0 (1.0) 99.9 ($<$0.1)
@@
We qualitatively analyze one-dimensional histograms of the typical shower shape variables. As shown in Fig 2, there is a close alignment between the results from GEANT4 data and generative models, which signifies that our generative model effectively captures the underlying patterns and distributions of the data. The consistent distribution patterns highlight the generative model’s fidelity in representing the essential features in the training dataset. Despite the observation of tail differences at the low fraction region for the positron in layer 1’s fraction histogram, the overall similarity between the performance of QVAE and DVAE models indicates our pipeline for $\beta_{\text{eff}}^{*}$ estimation and scaling is well behaved so that the quality of samples from the quantum device is on par with those from the block Gibbs method.
@@
To assess the practical deployment performance of the model’s energy conditioning, we generate samples from the trained model requesting a specific range of incident energy and histogram the observed energy. We generate the samples by classical and quantum approaches and compare the results with a GEANT4 test dataset under the same range of incident energy. As shown in Fig 3, for the positron cluster, the conditioning response of DVAE or QVAE can form a flat plateau with good accuracy in reproducing the defined selection of the input energy range. The pion cluster has a broadened response that is due to the nature of the unconstrained charged pion shower in the electromagnetic calorimeter. The shower shape histograms for both electron and pion clusters exhibit a noticeable similarity to the characteristics seen in the true data. This agreement indicates that our model has capability to capture and reproduce patterns from the first-principles simulation, suggesting its potential as a useful tool for calorimeter shower simulation.
@@
Figure 3: Comparison of energy spectra between CaloDVAE (classical) and CaloQVAE (quantum) models against GEANT4 test data for incident particle energies ranging from 45 to 60 GeV. The data encompasses both pion (top panel) and positron (bottom panel) clusters.
@@
The monitoring results of the DVAE sampling rate show efficient performance on our NVIDIA GeForce RTX 2080 Ti with 11264 MB memory. The processing time per sample for latent space sampling takes the most substantial portion ( $\sim$ 0.5 $\mathrm{ms}$), followed by data loading to the GPU ( $\sim$ 10 $\mathrm{ns}$), and passing through the decoder ($\sim$ 1 $\mathrm{\mu s}$). The DVAE model exhibits similar sampling rate with the CaloGAN results CaloGAN up to an order of magnitude. On the other hand, the quantum sampler achieves slightly faster sampling rate at 0.4 $\mathrm{ms}$ per sample, but impressively, the core QPU annealing rate is only 20 $\mathrm{\mu s}$ per sample, hinting at the raw speed of quantum processing. The extra QPU access time is mainly used for tasks like readout and handling delays, which have not been thoroughly optimized yet because these overheads are a much smaller fraction for more traditional quantum annealing applications. This suggests that with optimized engineering, the readout time could be optimized and quantum sampling could vastly outpace classical methods in efficiency for generative modeling, making it well-suited for real-time and large-scale applications.